OpenStudy (anonymous):

linearize the system about the equilibria or critical points. Identify the critical points, e.g. saddle, node, center,…, but do not sketch.

OpenStudy (anonymous):

\[\frac{ dx }{dt }=x-3y+2xy,\frac{ dy }{ dt }=4x-6y-xy\]

OpenStudy (anonymous):

I think I can use the Jacobian matrix

OpenStudy (michele_laino):

hint: we can write the Taylor series of the functions at the right side, up to the first order in x and y

OpenStudy (anonymous):

\[f(x)= f(x_{0})+f'(x_{0})(x-x_{0})+...\]

OpenStudy (anonymous):

We have not had a full lecture on taylor series yet, I am not familiar with it

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (michele_laino):

we have functions in 2 variables, so you have to write the Taylor series for functions of 2 variables

OpenStudy (anonymous):

I do not know how to do that.

OpenStudy (michele_laino):

I understand! Nevertheless the function: \[x - 3y + 2xy\] can be viewed as a function of 2 variables, namely: \[f\left( {x,y} \right) = x - 3y + 2xy\] and x=x(t), and y=y(t), similarly for the other function at the right side of the second differential equation

OpenStudy (anonymous):

Okay.

OpenStudy (anonymous):

\[g(x,y)=4x-6y-xy\]

OpenStudy (anonymous):

?

OpenStudy (michele_laino):

ok!

OpenStudy (michele_laino):

now, you have to apply the formula which gives the Taylor series for a function in 2 variables

OpenStudy (perl):

Consider the nonlinear system: \( \Large{ x ' = f(x,y) \\ y ' = g(x,y) } \) Near the critical points \( \large {(x_0,y_0) }\) satisfying \( \large{ f(x_0,y_0)=0 \\ g(x_0,y_0) =0 } \) we can approximate the non linear system by the linear system : $$\Large { \begin{pmatrix} x' \\ y' \\ \end{pmatrix} = \begin{pmatrix} \dfrac{\partial f}{\partial x}(x_o,y_o) & \dfrac{\partial f}{\partial y}(x_o,y_o) \\[1em] \dfrac{\partial g}{\partial x}(x_o,y_o) & \dfrac{\partial g}{\partial y}(x_o,y_o) \end{pmatrix} \begin{pmatrix} x \\ y \\ \end{pmatrix} } $$

OpenStudy (perl):

We can then analyze these linear systems to conclude what we can about the behavior of solutions near the critical points.

OpenStudy (perl):

we can use Michel's approach as well.

OpenStudy (perl):

its up to you

OpenStudy (anonymous):

I don't really know Michele's approach. I was trying to follow along with the Taylor series.

OpenStudy (anonymous):

Can we try your approach?

OpenStudy (perl):

ok i will continue

OpenStudy (perl):

we will use the post above as a template

OpenStudy (anonymous):

when I solved, I used the Jacobian matrix

OpenStudy (anonymous):

and okay!

OpenStudy (perl):

We are given \( \Large{ \frac{ dx }{dt }=x-3y+2xy,\\ \frac{ dy }{ dt }=4x-6y-xy } \) First task is to solve x' = 0 , y ' = 0 simultaneously

OpenStudy (perl):

Solve the system: x - 3y +2xy = 0 4x - 6y -xy = 0 clearly x=0, y=0 is a solution

OpenStudy (perl):

the other solution we can find by substitution

OpenStudy (perl):

Using the second equation 4x - 6y -xy = 0 4x - xy = 6y x ( 4 -y ) = 6y x = 6y / (4-y) substitute this into the first equation for x

OpenStudy (perl):

First equation: x - 3y +2xy = 0 substitute: 6y / (4-y) - 3y + 2 (6y / (4-y)) * y = 0

OpenStudy (perl):

$$ \large { x - 3y +2xy = 0 \\ \therefore \\ \frac{6y}{4-y} - 3y + 2 \cdot \frac{6y}{4-y} \cdot y =0 \\ \therefore \\ \color{red}{(4-y )} \cdot \left( \frac{6y}{4-y} - 3y + 2 \cdot \frac{6y}{4-y} \cdot y \right)=0 \cdot \color{red}{(4-y )} \\ \therefore \\ 6y - 3y(4-y) + 2 \cdot 6y \cdot y =0 \\ \therefore \\ 15y^2-6y = 0 \\ \therefore \\ y(15y-6) = 0 \\ \therefore \\y = 0, ~~y = \frac{6}{15}= \frac 25 } $$

OpenStudy (perl):

$$ \Large{ x' =f(x,y) = x-3y+2xy\\ y' =g(x,y) = 4x-6y-xy \\~\\ \begin{pmatrix} f_x & f_y\\ g_x & g_y \\ \end{pmatrix} = \begin{pmatrix} 1+2y & -3 + 2x\\ 4-y & -6 - x \\ \end{pmatrix} } $$

OpenStudy (perl):

now plug in the critical points

OpenStudy (anonymous):

\[\left[\begin{matrix}1 & -3 \\ 4 & -6\end{matrix}\right]\]

OpenStudy (anonymous):

thats for 0,0

OpenStudy (anonymous):

\[\left[\begin{matrix}1.8 & -3+2x \\ 3.6 & -6-x\end{matrix}\right]\]

OpenStudy (anonymous):

thats for (x, 2/5)

OpenStudy (perl):

actually the two critical points are (0,0) and (2/3, 2/5)

OpenStudy (anonymous):

\[\left[\begin{matrix}1.8 & -1 \\ 3.6& -6.667\end{matrix}\right]\]

OpenStudy (anonymous):

\[\left[\begin{matrix}1 & -3 \\ 3.6 & -6.667\end{matrix}\right]\]

OpenStudy (anonymous):

disregard the 1st one

OpenStudy (anonymous):

\[\left[\begin{matrix}1 & -3 \\ 4 & -6\end{matrix}\right]\]

OpenStudy (perl):

$$ \Large{ x' =f(x,y) = x-3y+2xy\\ y' =g(x,y) = 4x-6y-xy \\~\\ \begin{pmatrix} f_x & f_y\\ g_x & g_y \\ \end{pmatrix} = \begin{pmatrix} 1+2y & -3 + 2x\\ 4-y & -6 - x \\ \end{pmatrix} \\~\\ \text{The jacobian matrix associated with (0,0) is :} \\~\\ \begin{pmatrix} 1+2(0) & -3 + 2(0)\\ 4-0 & -6 - 0 \end{pmatrix} = \begin{pmatrix} 1 & -3\\ 4 & -6 \end{pmatrix} \\~\\ \text{The jacobian matrix associated with (2/3, 2/5) is :} \\~\\ \begin{pmatrix} 1+2(\frac25) & -3 + 2(\frac23)\\ 4-\frac25 & -6 - \frac23 \end{pmatrix} =\begin{pmatrix} \frac9 5 & -\frac 53\\ \frac{18}{5} & -\frac{20}{3} \end{pmatrix} } $$

OpenStudy (anonymous):

\[\left[\begin{matrix}1.8 & -1.667\\ 3.6& -6.667\end{matrix}\right]\]

OpenStudy (perl):

now we can use these matrices to make linear equations

OpenStudy (anonymous):

ex?

OpenStudy (anonymous):

x'=(9/5)x-(5/3)y y'=(18/5)x-(20/3)y

OpenStudy (perl):

$$ \Large{ \text{ near (0,0)}\\ \begin{pmatrix} x' \\ y' \end{pmatrix} =\begin{pmatrix} 1 & -3\\ 4 & -6 \end{pmatrix} \begin{pmatrix} x \\ y \end{pmatrix} \iff \begin{cases} x' = 1x -3y \\ y' = 4x - 6y \end{cases} \\ ~~\\~~\\ \\ \text{near (2/3, 2/5) } \\ \begin{pmatrix} x' \\ y' \end{pmatrix}=\begin{pmatrix} \frac9 5 & -\frac 53\\ \frac{18}{5} & -\frac{20}{3} \end{pmatrix} \begin{pmatrix} x \\ y \end{pmatrix} \iff \large \begin{cases} x' = \frac95x -\frac53y \\ y' = \frac{18}{5}x - \frac{20}{3}y \end{cases} } $$

OpenStudy (anonymous):

Ok, so now I have to identify the critical points?

OpenStudy (perl):

well now we can look at the eigenvalues

OpenStudy (perl):

Cases: Sources or Repellers: both eigenvalues have positive real part. Sinks or Attractors: both eigenvalues have negative real part. Saddles: one eigenvalues is positive and the other negative.

OpenStudy (anonymous):

for (0,0) --> sink ?

OpenStudy (perl):

we know the critical points are (0,0) and (2/3, 2/5)

OpenStudy (perl):

one moment, calculating eigen values

OpenStudy (perl):

for the first matrix associated with (0,0) i get eigen values -2, -3

OpenStudy (anonymous):

thats what I got

OpenStudy (perl):

so thats a sink :)

OpenStudy (anonymous):

yay!

OpenStudy (perl):

for the next matrix, i got eigenvalues Lambda = 1.019362877,-5.886029543

OpenStudy (perl):

Exactly: -73/30+1/30 sqrt(10729)],-73/30-1/30 sqrt(10729) This implies the critical point is a saddle

OpenStudy (anonymous):

thank you